### v1

# Inference-Aware Prompt Optimization for Aligning Black-Box Large Language Models

Abstract
Prompt optimization methods have demonstrated significant effectiveness in aligning black-box large language models (LLMs). In parallel, inference scaling strategies such as Best-of-N Sampling and Majority Voting have also proven to enhance alignment and performance by trading off computation. However, existing prompt optimization approaches are inference strategy agnostic; that is, they optimize prompts without regard to the inference strategy employed during deployment. This constitutes a significant methodological gap, as our empirical and theoretical analysis reveals a strong interdependence between these two paradigms. Moreover, we find that user preferences regarding trade-offs among multiple objectives and inference budgets substantially influence the choice of prompt and inference configuration. To address this gap, we introduce a unified novel framework named Iapo (Inference-Aware Prompt Optimization) that jointly optimizes the prompt and inference scale, while being aware of the inference budget and different task objectives. We then develop a fixed-budget training algorithm for Iapo, which we call Psst (Prompt Scaling via Sequential Trimming), and analyze finite-budget guarantees on error probability. Finally, we evaluate the effectiveness of Psst on six different tasks, including multi-objective text generation and reasoning, and demonstrate the critical role of incorporating inference-awareness when aligning black-box LLMs through prompt optimization.
Introduction
In recent years, most state-of-the-art large language models (LLMs) are accessible only through black-box APIs. Traditional alignment methods that require access to model weights or logits are therefore infeasible. To address this issue, prompt optimization-based alignment methods have garnered interest (Chang et al. 2024). These methods typically enhance input prompts by rewording or appending additional instructions to better align the models’ outputs with a task’s objectives. Another broadly applicable alignment strategy for black-box models is scaling inference computations using strategies such as Best-of-N Sampling or Majority Voting. These inference scaling methods generate multiple candidate responses for the same query and select the final response via ranking or voting mechanisms (Wang et al. 2022; Krishna et al. 2022; Gui, Gârbacea, and Veitch 2024; Yue et al. 2025).
Although existing prompt optimization techniques have achieved substantial success, they are typically agnostic to how model outputs are aggregated or sampled, overlooking the impact of such inference methods. Our initial empirical investigation reveals that the performance of optimized prompts is highly sensitive to the choice of inference-scaling approach. Furthermore, our theoretical analysis reveals that decoupling prompt optimization from inference can lead to misalignment. Finally, we observe that optimal alignment requires careful consideration of user-specific preferences regarding the trade-offs among multiple objectives, as well as the computational resources they are willing to expend. These findings expose a critical gap in current methods: the absence of a unified framework that simultaneously accounts for prompt optimization, inference-scaling strategies, user preferences, and computational resource constraints.
To bridge this gap, we introduce Iapo (Inference-Aware Prompt Optimization), a novel prompt optimization framework designed explicitly to produce aligned responses from inference-scaled black-box LLMs. Iapo simultaneously optimizes prompt design and inference scaling strategies while considering different task objectives and computational budgets. We formulate the task of identifying an optimal policy for the Iapo framework as a contextual best-arm identification (BAI) problem. To efficiently solve this, we propose a fixed-budget training algorithm named Psst (Prompt Scaling via Sequential Trimming). Additionally, we introduce a warm-up heuristic that further improves performance within the training budget.
We begin our analysis by deriving theoretical finite-budget guarantees on the error probability of Psst. Next, we empirically demonstrate the effectiveness of Psst for learning Iapo policies across six diverse tasks, including multi-objective text generation, mathematical reasoning, and commonsense reasoning benchmarks. Additionally, our analysis shows that ignoring inference scaling during prompt optimization can lead to substantial misalignment, highlighting the critical role of inference-awareness in aligning black-box LLMs.
Related Work
Over the years, considerable effort has been devoted to aligning large language models (LLMs) with human expectations in downstream tasks (Minaee et al. 2024). Many widely adopted alignment approaches—such as Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF), and Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert 2025)—require access to model weights. This limitation has motivated increasing interest in black-box alignment methods such as prompt optimization, which can align black-box models through input manipulation alone (Zhou et al. 2022; Ouyang et al. 2022; Chang et al. 2024). Prompt optimization has demonstrated strong performance in both single-objective (Cheng et al. 2023; Trivedi et al. 2025) and multi-objective (Jafari et al. 2024; Zhao et al. 2025) settings. However, these methods are agnostic of the inference strategy used during deployment, potentially leading to suboptimal performance. In contrast, our approach explicitly addresses the interdependence of inference-time strategy and prompt optimization.
Recently, Shi et al. framed prompt optimization as a fixed-budget best-arm identification (BAI) problem. While effective under limited evaluation budgets, the method remains inference agnostic and was only explored in single-objective settings. Our work builds on this foundation in two key ways: (1) we introduce a contextual formulation that models user preferences over multiple objectives and associated computational costs; and (2) we incorporate inference-awareness to ensure alignment with the actual inference strategy. To learn an optimal policy, we introduce a fixed-budget contextual BAI algorithm, Psst, inspired by Sequential Halving (SH) (Karnin, Koren, and Somekh 2013). While SH was originally developed for the pure bandit setting, the Iapo framework features both inter-context full-information feedback and intra-context semi-bandit feedback. Psst leverages these structural properties to achieve more efficient optimization, extending beyond what standard SH can accommodate.
Another relevant line of work focuses on inference-time alignment, where model outputs are improved during inference without modifying model parameters. Some of these methods, such as GenARM and DEAL (Xu et al. 2024; Huang et al. 2024), require access to model logits, limiting their applicability in black-box settings. In contrast, Best-of-N sampling (BoN) and Majority Voting (MV) methods operate purely on model outputs and have shown strong empirical gains by generating multiple candidates and selecting the best one (OpenAI 2024; Yue et al. 2025; Wang et al. 2022; Krishna et al. 2022). However, these approaches introduce a non-trivial computational cost, and to our knowledge, none of them explicitly optimize the trade-off between computational budget and output quality. Further, our preliminary experiments show that such inference-scaling strategies interact non-trivially with prompt design: prompts optimized for single-shot decoding may yield suboptimal performance under BoN or MV, and vice versa. This necessitates an inference-aware prompt optimization framework.
Finally, some white-box methods have recently integrated inference-awareness into the training process. For example, Chow et al. (2025) proposed an inference-aware fine-tuning procedure that explicitly optimizes for exploration–exploitation trade-offs under BoN. Similarly, BOND (Sessa et al. 2024) and BonBon (Gui, Gârbacea, and Veitch 2024) aim to distill BoN policies into a single-pass generation process through supervised fine-tuning. While these approaches avoid the cost of sampling at inference time, they require full access to model parameters and do not generalize beyond BoN-style strategies. In contrast, our method is complementary to inference-aware fine-tuning designed for black-box LLM.
Inference–Aware Prompt Optimization
In this section, we first formalize the problem setup and introduce the Iapo framework. Next, we present an empirical example that highlights the need for inference-aware optimization. Building on these observations, we then establish theoretical conditions under which Iapo is necessary compared to disjoint optimization.
Problem Formulation
Let $\mathcal{X}$ be the distribution of user queries and $\mathcal{P}$ a finite prompt set. A pair $(x\in\mathcal{X},p\in\mathcal{P})$ is submitted to a frozen black-box LLM, which, under fixed decoding hyperparameters, generates $N\in\{1,\dots,N_{\max}\}$ i.i.d. completions $\mathbf{y}_{1:N}=(y_{1},\dots,y_{N})$. $K$ bounded (potentially vector) objectives score each completion $O_{k}:\mathcal{X}\times\mathcal{Y}\to[o_{k}^{\min},o_{k}^{\max}]$ (e.g. helpfulness, harmlessness, exact-match). We also define the cost of producing a response as $\mathrm{Cost}(x,y_{i})$, a composite function that takes into account various computational factors such as token count, time, and energy. We add it as a (K+1)(K{+}1)-st objective $O_{k+1}=-\mathrm{Cost}(x,y_{i})$. An external entity supplies a context $c=(w_{1},\dots,w_{K+1})\in\mathcal{C}$, where every $w_{k}$ is chosen from a finite discrete domain. Given the above setup, we formalize the inference strategy as follows.
Best-of-N (BoN).
BoN returns the largest weighted utility:
$R_{x}^{\textsc{BoN}}(c,p,N)=\underbrace{\max_{i\leq N}\sum_{k=1}^{K}w_{k}\,o_{k}\bigl{(}x,y_{i}\bigr{)}}_{\text{task reward}}+\underbrace{w_{k+1}\sum_{i=1}^{N}o_{k+1}(x,y_{i})}_{\text{inference cost}}.$ (1)
Majority Voting (MV).
For query $x$, the pair $(p,N)$ yields i.i.d. completions $y_{1:N}$ and extracted answers $\ell_{i}=h(x,y_{i})$. For each distinct answer $s$, define the vote count $n_{s}=\sum_{i=1}^{N}\mathbf{1}[\ell_{i}=s]$, the maximum $n^{\star}=\max_{s}n_{s}$, and the tie multiplicity $t=\sum_{s}\mathbf{1}[n_{s}=n^{\star}]$. MV predicts uniformly at random among the $t$ maximizers. With gold answer $a(x)$ and the success credit defined as $o_{1}(x,p,N)=\frac{\mathbf{1}[\,n_{a(x)}=n^{\star}\,]}{t}$ we define MV utility as:
$R_{x}^{\textsc{MV}}(c,p,N)=\underbrace{w_{1}\,o_{1}(x,p,N)}_{\text{task reward}}\;+\;\underbrace{w_{2}\sum_{i=1}^{N}o_{2}(x,y_{i},c)}_{\text{inference cost}}.$ (2)
Remark.
A mixed strategy arises when different objectives require different aggregation rules, e.g., applying MV for binary correctness and BoN for stylistic quality in reasoning tasks. It is trivial to define it on the basis of the above.
(a) MATH, Majority Vote
(b) Helpful-Harmless, Best-of-N (Prompt-A)
(c) Helpful-Harmless, Best-of-N (Prompt-B)
Figure 1: Prompt–Inference Interdependence. (a) Accuracy under Majority Voting with LLaMA-3.3-70B-Instruct, showing prompt dominance shifts with budget (shaded). (b, c) Cost-adjusted reward under Best-of-N decoding. Prompt and inference scales vary with user-defined trade-offs.
Iapo Framework
Let an inference configuration be a tuple $\theta\in\Theta$ (e.g. temperature, top-$p$, max token). Then we define a set of arms $\mathcal{A}$ in Iapo as: $a\;=\;(p,\theta,N)\;\in\;\mathcal{A}:=\mathcal{P}\times\Theta\times\{1,\dots,N_{\max}\}.$
Thus, each arm fixes the prompt, the decoding hyperparameter, and the number of sampled completions. However, throughout the text, we fold the inference configuration into the prompt $p$ and write $a=(p,N)$. Finally, an Iapo policy is defined as a mapping $\pi:\mathcal{C}\;\rightarrow\;\mathcal{A}$ that selects an arm after observing a context $c$.
Given a dataset $\mathcal{X}$, context $c\in\mathcal{C}$, and aggregator $\alpha\in\{\textsc{BoN},\textsc{MV}\}$ the expected utility of arm $a$, i.e., the context-action value function or $Q$-function is defined as:
$Q^{\alpha}(c,a):=\mathbb{E}_{x\sim\mathcal{X}}\bigl{[}R_{x}^{\alpha}(c,a)\bigr{]}.\quad$ (3)
Note that $R_{x}^{\alpha}(c,a)$ is a random variable. Now, let the context-optimal arm is $a^{\star}(c)=\arg\max_{a}Q^{\alpha}(a,c)$; hence the optimal Iapo policy is defined as: $\pi^{\star}(c)\;=\;a^{\star}(c),\forall\,c\in\mathcal{C}$.
In this paper, we adopt a train-then-deploy setup to learn the optimal IAPO policy. Given a total completion budget of $T$, the learner may adaptively select arms $a_{t}=(p_{t},N_{t})\in\mathcal{A}$ and query $x_{t}\sim\mathcal{X}$, then observe the full raw reward vector $\mathbf{m}_{t}\in\mathbb{R}^{K+1}$ for all completions. This process may continue until the budget is exhausted ($\sum_{t}N_{t}=T$). After spending the entire budget, the learner returns a deployment policy $\pi_{T}$. The performance of this policy is evaluated by the Average Contextual Return:
$\text{ACR}(\pi_{T})\;=\;\mathbb{E}_{c\sim\mathcal{C}}\bigl{[}Q^{\alpha}(c,a)\bigr{]},$ (4)
The goal of a learning algorithm is to return a deployment policy $\pi_{T}$ for a fixed pull budget $T$ that maximizes the ACR.
Motivating Case Study
To illustrate the limitations of inference-agnostic prompt optimization—and to motivate the joint treatment formalized above—we conducted two diagnostic experiments with Llama-3.3-70B-Instruct (Grattafiori et al. 2024) strictly treated as a black-box API. The results are summarized in Figure 1.
(a) Majority Voting on MATH.  We evaluate three manually designed prompts on the MATH benchmark (Hendrycks et al. 2021) under Majority Voting with $N\in\{1,\dots,16\}$. Accuracy is plotted against total decoding cost, averaged over 300 queries (see the Appendix for details). Two key observations emerge. First, prompt preference shifts with compute budget: the green prompt performs best at low budget, but is eventually surpassed by the blue prompt as Majority Voting becomes more effective. Second, inference-agnostic optimization can be short-sighted: selecting a prompt based solely on single-shot (N=1N{=}1) accuracy would favor the green prompt, overlooking the fact that the blue prompt is strictly superior for any user willing to allocate more compute.
To see how the green and blue trend can emerge, consider the following example. Suppose in a reasoning task with MV, for Prompt 1 we have $40\%$ in Query 1, $90\%$ in Query 2, and for Prompt 2, $62\%$ (both queries). Single-shot averages favor A ($0.65$ vs. $0.62$), but under MV with $N=10$, A drops to $\approx 0.63$ while B improves to $\approx 0.77$.
(b,c) Best-of-$\boldsymbol{N}$ on Helpful–Harmless. We evaluate two prompts (A and B) on the Helpful–Harmless benchmark (Bai et al. 2022) using Best-of-$N$ decoding for $N\leq 24$. Each curve corresponds to a different user-defined trade-off between helpfulness and harmlessness, plotting the cost-adjusted reward averaged over 1000 queries (see the Appendix for details). The optimal choice of prompt (A vs. B) and sampling budget (N) is highly sensitive to these preferences. For example, the prompt A is strictly preferred when helpfulness is weighted more heavily.
Having established the need for inference-aware optimization, we now examine the precise conditions under which joint optimization becomes essential. We start by establishing the Inference-Agnostic (IA) utility:
Proposition 1 (Inference-Agnostic Utility).
Inference-agnostic prompt-optimization methods optimize cost-unaware arithmetic mean utility.
$R_{x}^{\textsc{IA}}(c,a=(p,N))=\frac{1}{N}\textstyle\sum_{i=1}^{N}\sum_{k=1}^{K}w_{k}o_{k}(x,y_{i}).$ (5)
Now we show under what conditions the IA policy remains optimal or an optimal policy can be trivially recovered from the IA $Q$-function.
Proposition 2 (Inference-Agnostic Optimality).
The Inference-Agnostic prompt-optimization policy remains optimal under linear transformation of $R_{x}^{\textsc{IA}}(c,a)$, that is, $kR_{x}^{\textsc{IA}}(c,a),k>0$ and an optimal policy can be recovered trivially from $Q$-function under affine transformation:
$Q^{AF}(c,a):=\mathbb{E}_{x\sim\mathcal{X}}\bigl{[}aR_{x}^{IA}(c,a)+b\bigr{]}=kQ^{IA}(c,a)+b.\quad$
Figure 2: Expected utility ($w_{k+1}=0$) for MV (left) and BoN (right). MV shows a sharp performance drop when the correctness probability drops below 0.5, whereas BoN is strictly concave.
The above also highlights that affine aggregation significantly simplifies inference-aware optimization. For instance, in a regression task where the aggregated prediction is the mean of multiple numeric predictions and the reward is defined by the mean squared error (MSE), in some cases can become an affine transformation of IA, eliminating the need to simulate inference scaling during training. However, common inference scaling strategies like BoN and MV generally do not admit such affine formulations. While they can sometimes be expressed as non-affine transformations of the IA—such as in the Bernoulli case with large $N$, where $R_{x}^{\textsc{IA}}(c,a)\approx p$ (Figure 2)—these are special cases. Hence, trying to determine the prompt based on $Q^{IA}$ for BoN or MV will result in misalignment. This motivates the next section, where we develop a training method that handles the general Iapo setting beyond the affine regime.
Prompt Scaling via Sequential Trimming
In this section, we propose a fixed-budget arm elimination-based strategy for training policy $\pi_{T}$, called Psst (Prompt Scaling via Sequential Trimming). We then provide a theoretical analysis that establishes error guarantees for Psst under a finite inference budget. Finally, we introduce a practical approximation heuristic that improves computational efficiency without significantly compromising performance in many practical settings.
Our focus on the fixed inference budget setting is motivated by the fact that training cost is often the main bottleneck in real-world applications. Moreover, Psst is designed to operate in a batched-exploration mode, which further reduces costs since many black-box APIs offer significant discounts for batched inference compared to individual calls. Importantly, Psst is also hyper-parameter-free, requiring no additional tuning.
Classical arm–elimination methods such as Sequential Elimination (Even-Dar, Mannor, and Mansour 2006) and Sequential Halving (Karnin, Koren, and Somekh 2013) follow a simple recipe: (i) split the elimination process into multiple rounds; (ii) in each round, allocate the round budget across the surviving arms; and (iii) trim a subset of arms at the end of the round based on their estimates. However, Iapo departs from pure BAI settings in the following key ways:
  • Asymmetric pull cost. When arm $(p,N)$ is pulled during training, it uses $N$ training budget.
  • Cross‑context reuse. One pull of $(p,N)$ on query $x$ yields the completion set $\mathbf{y}_{1:N}$ and objective vector set $\mathbf{o}_{1:N}$ that can can be used to estimate $R^{\alpha}_{x}(c,p,N)$ for all $c\in\mathcal{C}$.
  • Nested sample reuse across inference scales. Pulling a larger scale subsumes smaller ones: a pull of $(p,N_{i})$ produces $\big{\lfloor}N_{i}/N_{j}\big{\rfloor}$ i.i.d. block samples for arm $(p,N_{j})$ by partitioning the $N_{i}$ draws into disjoint groups of size $N_{j}$ (e.g., to recompute BoN/MV on each block).
A key consequence is that, for a prompt, the largest surviving scale drives the budget. Let $N^{(r)}_{\max}(p)=\max\{\,N:\ (p,N)\ \text{survives at the start of round }r\,\}.$ If we allocate $K$ pulls (blocks) to $(p,N^{(r)}_{\max}(p))$ in round $r$, then every surviving arm $(p,N)$ with $N\leq N^{(r)}_{\max}(p)$ automatically receives at least $K$ effective samples by block reuse. Thus, an effective arm elimination strategy should exploit both (i) cross‑scale reuse within prompts and (ii) cross‑context reuse when scoring, while being aware of asymmetric cost.
Algorithm 1 Prompt Scaling via Sequential Trimming
1:Context set $\mathcal{C}$, prompt set $\mathcal{P}$, scale set $\mathcal{N}$, Scaling strategy $\alpha$, Query Dataset $\mathcal{X}$, total pull budget $T$;
2:for all $(c,a)\,\in\,\mathcal{C}\times\mathcal{A}$ do
3:  $\text{F}_{c,a}\leftarrow\textbf{true}$
4:end for
[…]
6:for $r=1$ to $R$ do
7:  $n_{r}\leftarrow\bigl{\lfloor}T/R\bigr{\rfloor}$
8:  $\lambda^{(r)}\leftarrow\textsc{Allocate}\bigl{(}\text{F},n_{r}\bigr{)}$
9:  $\mathcal{B}\leftarrow\{\}$
10:  for $(a,n_{r})\in\lambda^{(r)}$ do
11:   for $i=1...n_{r}$ do
12:     Sample $x\sim\mathcal{X}$
13:     $\mathcal{B}\leftarrow\mathcal{B}\ \cup(a,x)$
14:   end for
15:  end for
16:  $\mathcal{D}\leftarrow\textsc{Batch-Query}\bigl{(}\mathcal{B})$
17:  $Q^{\alpha}_{(r)}\leftarrow\textsc{Estimate-Q}(\mathcal{D})$
18:  for all $c\in\mathcal{C}$ do
19:   $\mathcal{A}_{c}^{(r)}\leftarrow\{a:\text{F}_{c,a}=\textbf{true}\}$
20:   Rank $\mathcal{A}_{c}^{(r)}$ by $Q^{\alpha}_{(r)}(c,a)$
21:   Remove bottom $\lceil|\mathcal{A}_{c}^{(r)}|/2\rceil$ arms // i.e. update $\mathbf{F}$
22:  end for
23:end for
24:return $\{a_{c}^{\star}\}_{c\in\mathcal{C}}$ // one survivor per context
Round Structure.
Algorithm 1 proceeds in $R=\lceil\log_{2}|\mathcal{A}|\rceil$ rounds, and tracks per context active arm using the flag $\mathbf{F}$. Each round is allocated an equal pull budget of $n_{r}=\lfloor T/R\rfloor$. An allocation routine, Allocate(F, $n_{r}$), divides this budget among the current set of unique active arms, aggregated across all contexts. Based on this allocation, a batch of inference calls is issued to the target LLM. The resulting completions are scored using a reward function or verifier and stored in the dataset $\mathcal{D}$. The $Q$-values are then estimated from the collected data. Within each context, arms are ranked and the worst performing half are eliminated. After all rounds are completed, the algorithm returns a single final arm for each context.
Structure-Aware Allocation Policy.
The allocation policy is designed with cross-context and cross-scale information sharing in mind. Specifically, let $\mathcal{A}^{(r)}$ denote the set of unique active arms in round $r$, aggregated across all contexts. For each prompt $p$, define
$N^{(r)}_{p,\max}=\max\{N\mid(p,N)\in\mathcal{A}^{(r)}\}$
as the maximum inference scale for prompt $p$ among the active arms. Then, Psst allocates the budget to each arm according to the following scheme:
$\lambda(a)=\begin{cases}(\lfloor\frac{n_{r}N^{(r)}_{p,max}}{M}\rfloor&\text{if }a=(p,N_{p}^{\max})\in\mathcal{A}^{(r)},\\ 0&\text{otherwise},\end{cases}$ (6)
where $M=\sum_{p:(p,N_{p}^{\max})\in\mathcal{A}^{(r)}}N^{(r)}_{p,\max}$ is the total cost of sampling all such maximal arms once. This policy maintains uniform coverage over prompts while respecting cost asymmetries and ensures that the maximum scale of every prompt has an equal number of samples.
We now derive error bounds [^1] Proof in the appendix for Psst under the allocation policies described above.
Theorem 1 (Error of PSST).
Let $R=\lceil\log_{2}|\mathcal{A}|\rceil$ be the number of trimming rounds, and $o_{k}^{\min},o_{k}^{\max}]=[-1,1]$ and define the cost–gap complexity
$H_{1}^{c}=\max_{(c,a_{i})\neq a^{c}_{1}}\frac{\bar{N}_{max}}{\Delta_{c,a^{i}}^{\,2}},\quad H_{1}=\max_{c}H_{1}^{c}.$
$H_{2}^{c}=\max_{(c,a_{i})\neq a^{c}_{1}}\frac{i\bar{N}_{max}}{\Delta_{c,a^{i}}^{\,2}},\quad H_{2}=\max_{c}H_{2}^{c}.$
$\text{where},\ \Delta_{c,a^{i}}=Q^{\alpha}_{c,a^{1}}-Q^{\alpha}_{c,a^{i}}$, arms are indexed based on ascending order of $Q^{\alpha}_{c,a}$ under that context and $\bar{N}_{max}=\frac{a_{1}(N)+N_{max}}{2}$. Running PSST with the structure-aware allocation of for a total prompt complication $T$ returns the optimal arm in every context with probability at least
$1-3|\mathcal{C}|R\exp\!\Big{(}-\tfrac{T}{min(2|\mathcal{P}|H_{1},8|C|H_{2})R}\Big{)}.$
Equivalently, to ensure failure probability at most $\delta$ it suffices to choose
$T=O\Big{(}min(|\mathcal{P}|H_{1},|\mathcal{C}|H_{2})R\,\log\!\Bigl{(}\tfrac{|\mathcal{C}|R}{\delta}\Bigr{)}\Big{)}.$
Note that applying Sequential-Halving without leveraging the structure of Iapo—specifically, without any form of information sharing across scales or contexts—incurs a sample complexity of $O(|\mathcal{C}|N_{\max})$ higher.
Remark:
While we describe the algorithm as where we use a new set of data in each round $\mathcal{D}$, it has been shown that in similar halving-style algorithms (Fabiano and Cazenave 2021), data accumulating all past observations—known as stockpiling—can improve the complexity of $T$ by reducing the outer $R$-factor, and is recommended to use with Psst.
Top-$\boldsymbol{K}$ Screening.
To further reduce the budget requirement of Psst, we introduce Top-$\boldsymbol{K}$ Screening, a practical heuristic that executes a short, uniform prompt screening at unit scale to trim obviously suboptimal prompts before running full PSST. Top-$\boldsymbol{K}$ Screening takes a budget fraction $T_{0}=\lfloor\rho T\rfloor$ ($\rho\in(0,1)$) from Psst. With scale restriction of N=1N{=}1, the budget is allocated uniformly across prompts: each $p\in\mathcal{P}$ receives $\big{\lfloor}T_{0}/|\mathcal{P}|\big{\rfloor}$ i.i.d. samples. Based on this data, $Q^{\alpha}(c,p,1)$ is estimated $\forall c\in\mathcal{C},p\in\mathcal{P}$. For each context $c$, we retain the $K$ best prompts $\mathcal{P}^{(0)}_{c}=\mathrm{Top}\text{-}K\{\,\widehat{Q}^{\alpha}(c,p,1):p\in\mathcal{P}\,\}$ and discard the rest. The subsequent PSST run is then restricted to the reduced arm sets $\mathcal{A}^{(1)}_{c}=\{(p,N):p\in\mathcal{P}^{(0)}_{c},\ N\in\mathcal{N}\}$ for each $c$, and uses the remaining budget $T^{\prime}=T-T_{0}$. In the next section, we demonstrate that the screening strategy can significantly improve performance in low training budget settings without compromising quality for practical tasks. However, theoretical guarantees comparable to those of full Psst cannot be established; counterexample tasks can be carefully constructed within Iapo framework, where Top-K screening will behave suboptimally for any $K<|\mathcal{P}|$.
Empirical Evaluation
Figure 3: Comparison between exploration strategies across six datasets.
In this section, we empirically evaluate the effectiveness of Psst and highlight the importance of inference-aware prompt optimization (Iapo). Our evaluation has two primary objectives:
  • To demonstrate that Psst and the Top-K Screening heuristic are highly effective at learning policy $\pi_{T}$.
  • To show that Iapo improves the average cost-adjusted reward (ACR) compared to inference strategy agnostic optimization.
Baselines.
We compare Psst and Top-K Screening with several baselines. We denote Top-K Screening with $K=1,K=4$,and $K=8$ as Psst+K1+K1, Psst+K4+K4, and Psst+K8+K8 respectively. For these heuristics, we fix $\rho=0.2$, which was found to perform best across all datasets. Full Psst is parameter-free and does not require any tuning. In our first set of experiments, we compare our proposed methods against several standard exploration strategies:
  • Uniform: Uniformly explores all arms in one batch and selects the best arm at the end.
  • $\boldsymbol{\epsilon}$-greedy: Samples a random context at each step and selects the best arm with probability $1-\epsilon$. We set $\epsilon=0.15$, which yielded the best performance across datasets.
  • Softmax: Samples arms according to a softmax distribution over estimated $Q$ values.
  • UCB: At each turn, selects the arm with the highest optimistic $Q$ estimate. The exploration constant $0.1$ after tuning.
Note that all baseline methods share information across contexts and inference scales; however, none of them are designed to exploit Iapo structure, i.e., they are structure-agnostic.
In the second set of experiments, we consider the well-known contextual variant of TRIPLE-SH (Shi et al. 2024) method, which optimizes prompt selection as a pure best-arm identification (BAI) problem. However, it does not optimize the inference scale. Therefore, we include two variants:
  • TRIPLE (N = 1): Only performs prompt optimization with single-sample inference.
  • TRIPLE (N = Random): Optimizes the prompts while randomly assigning $N$ for each query.
These baselines help isolate the benefits of jointly optimizing prompts and inference scale. Further, Psst+K1+K1 is particularly interesting in this experiment, as it approximates a two-stage disjoint optimization: it first selects a context-specific single-shot prompt using a cost-aware objective, and then tunes the inference scale. The Psst+K4+K4 and Psst+K8+K8 heuristics represent intermediate strategies between disjoint and fully joint optimization.
Note that all hyperparameter sweep results are in the supplementary material; we report results with the best setting found across all six datasets.
Environments.
Environments $\alpha$ $\lvert\mathcal{P}\rvert$ $N_{\max}$ $o_{k}^{max}$ $|\mathcal{X}|$ $|\mathcal{C}|$
Synth–Bernoulli MV 32 32 1.0 520 3
MATH MV 25 32 1.0 316 3
CommonsenseQA MV 48 32 1.0 1500 3
Synth–Categorical BoN 32 32 4.0 512 27
Helpful–Harmless BoN 20 32 1.0 1355 27
Summarization BoN 20 32 1.0 1201 27
Table 1: Environment summary.
Figure 4: Effectiveness of inference-aware optimization across six datasets.
We evaluated inference-aware optimization across a total of six environments. Key details are provided in Table 1. Environments one and four are synthetically constructed to mimic Iapo tasks, where prompt-query pair score distributions $o_{i}(c,P,1)$ are modeled using categorical distributions. We introduce them to validate some of the theoretical findings. The remaining four environments are based on widely-used real-world datasets. Among these, MATH (Hendrycks et al. 2021) and CommonsenseQA (Talmor et al. 2018) are used to evaluate reasoning tasks under Majority Voting (MV), while Helpful-Harmless (Bai et al. 2022) and Summarization (Stiennon et al. 2020) are chosen for Best-of-N (BoN) evaluation.
For the MV tasks, the task objective is defined as an exact match with the correct answer. All three BoN tasks are bi-objective, and we use publicly available reward models from previous multi-objective LLM alignment studies to score completions (see appendix for links). The cost objective in all six tasks is defined to be proportional to the average number of tokens per response. For context specification, MV tasks include a budget regime \$low, mid, high\$, while BoN tasks include both the budget and the bi-objective weights, which range from 0.1 to 0.9 for each objective. For example, in the helpful–harmless task, a context might be represented as $\{\text{helpful}:0.3,\text{harmless}:0.7,\text{budget}:\text{high}(1.0)\}$. Further details, including all prompts, are provided in the supplementary material.
To construct the environments, we first generated a set of instruction prompts for each task using ChatGPT-O3. We then generated 128 responses for each prompt–query pair and estimated the score distribution using a categorical model. All completions were produced using the LLaMA-3.3-70B-Instruct, a widely used open-source model (Meta_AI 2024), which we treat as a black-box throughout our experiments. Generation was carried out with vLLM(Kwon et al. 2023) on a cluster of 8 A100 GPUs, totaling approximately 2,000 GPU-hours. Once the environments are constructed, all experiments can be run via a standard CPU quickly. We will publish the environments and code with the paper, enabling full reproducibility without any substantial computational resources.
Evaluation Protocol.
All reported curves are averages over 200 independent runs. For synthetic environments, we instantiate 200 independent environments and report the average performance across them. For the remaining four environments, each run reshuffles the dataset, performs an $80/20$ train–test split, and trains the policy on the training set. In all six environments, we evaluate ACR on the test set using $10,000$ samples. Performance for each budget is the mean across the 200 runs, with standard error of the mean (SEM) error bars. Statistical significance is assessed using the Wilcoxon paired two-sided test with alpha $0.05$, and we indicate when differences are significant in the discussion. The full set of results is in the appendix.
Comparison of Exploration Strategies (Fig. 3).
Psst and the Top-K screening heuristic consistently outperform all baselines. Across all six domains, where the per-context action spaces are large ($|\mathcal{P}|N_{\max}\in[640,1536]$), UCB, softmax, and $\varepsilon$-greedy methods struggle to explore effectively. Among the baselines, UCB performs comparably in some domains after $T=20\text{K}$, but only with extensive hyperparameter tuning. Furthermore, these baselines are fully sequential and cannot leverage the cost and computational efficiency benefits of batch exploration. Full Psst attains the best final performance across four settings, while Psst+KXK\!X typically reaches strong policies faster, matching or exceeding Psst on three of the four real‑data tasks when the budget is small. Under aggressive pruning (small $K$), however, the heuristic becomes suboptimal—most notably on summarization and on the synthetic benchmarks—suggesting that Psst+KXK\!X is attractive under tight budgets, whereas full Psst is preferable for critical tasks such as long‑horizon, high‑frequency deployment. Finally, the statistical test also validates that Psst, along with Top-K screening, significantly outperforms baselines in all six datasets and under nearly all budgets. These findings indicate that our approach reliably discovers well-aligned solutions using as few as $5\text{K}$ to $20\text{K}$ inference calls in practical settings.
Importance of Inference-Awareness (Fig. 4).
We examine the role of inference awareness in prompt optimization. Across all six datasets, Iapo methods markedly outperform the inference-agnostic methods, demonstrating the gains achievable when jointly optimizing the prompt and inference scale. $\mathit{TRIPLE}(N=1)$ fails as it does not leverage inference scaling. On the other hand, $\mathit{TRIPLE}(N=Random)$ fails because it does not optimize the scaling for different contexts. The screening variant PSST+K1+K1—which effectively approximates a near‑decoupled (prompt‑only) procedure—fails to reach the optimum in most cases, performing competitively only on CommonsenseQA and showing pronounced underperformance on summarization. This is because it gets stuck with deceiving prompts that fail to scale compared to prompts that may not perform well under single-shot but improve significantly under scaling. These findings underscore the essential role of Iapo in aligning black‑box LLMs and the pitfalls of disjoint optimization. Overall, Iapo outperforms disjoint optimization by up to 25% and prompt-only optimization by up to 50%.
Conclusions and Future Work
We present an inference-aware prompt optimization (Iapo) framework for aligning black-box LLMs, emphasizing that prompts and deployment-time inference scaling strategy are tightly coupled and should be optimized jointly. Our proposed Psst and Top-K Screening heuristic demonstrate consistent improvements over strong baselines across six different settings. Looking ahead, we plan to explore richer inference-scaling policies (e.g., adaptive BoN/MV schedules, stopping rules, and tree search). We also aim to extend the framework to multi-objective alignment with explicit cost/latency constraints and to study long-horizon deployments under distribution shift.
Acknowledgments
This research was supported in part by the U.S. Army DEVCOM Analysis Center (DAC) under contract number W911QX23D0009, by the National Science Foundation grants 2205153, 2321786, and 2416460, and by Schmidt Sciences under the AI Safety Science program.
Appendix A
Proof of Theorem 1
Theorem 2 (Error of PSST).
Let $R=\lceil\log_{2}|\mathcal{A}|\rceil$ be the number of trimming rounds, and $o_{k}^{\min},o_{k}^{\max}]=[-1,1]$ and define the cost–gap complexity
$H_{1}^{c}=\max_{(c,a_{i})\neq a^{c}_{1}}\frac{\bar{N}_{max}}{\Delta_{c,a^{i}}^{\,2}},\quad H_{1}=\max_{c}H_{1}^{c}.$
$H_{2}^{c}=\max_{(c,a_{i})\neq a^{c}_{1}}\frac{i\bar{N}_{max}}{\Delta_{c,a^{i}}^{\,2}},\quad H_{2}=\max_{c}H_{2}^{c}.$
$\text{where},\ \Delta_{c,a^{i}}=Q^{\alpha}_{c,a^{1}}-Q^{\alpha}_{c,a^{i}}$, arms are indexed based on ascending order of $Q^{\alpha}_{c,a}$ under that context and $\bar{N}_{max}=\frac{a_{1}(N)+N_{max}}{2}$. Running PSST with the structure-aware allocation of for a total prompt complication $T$ returns the optimal arm in every context with probability at least
$1-3|\mathcal{C}|R\exp\!\Big{(}-\tfrac{T}{min(2|\mathcal{P}|H_{1},8|C|H_{2})R}\Big{)}.$
Equivalently, to ensure failure probability at most $\delta$ it suffices to choose
$T=O\Big{(}min(|\mathcal{P}|H_{1},|\mathcal{C}|H_{2})R\,\log\!\Bigl{(}\tfrac{|\mathcal{C}|R}{\delta}\Bigr{)}\Big{)}.$
Lemma 1.
The probability that the best arm under context $c$ is eliminated from context $c$ on round $r$ is at most
$2\,\exp\!\Big{(}-\tfrac{T}{2|\mathcal{P}|H_{1}^{c}R}\Big{)}$
Proof.
Assume that the best arm was not eliminated before round $r$. Then due to Hoeffding’s inequality for any arm $a^{i}\in\mathcal{A}^{(r)}_{c}$,
$\Pr\!\big{[}\hat{Q}^{\alpha,(r)}_{c,a_{1}}<\hat{Q}^{\alpha,(r)}_{c,a_{i}}\big{]}\;\leq\;\exp\!\left(-\frac{1}{2}harmonic(t_{r_{1}},t_{r_{i}})\,\Delta_{c,a_{i}}^{2}\right).$
Here, $t_{r}$ is the number of samples that were used to estimate the $Q$ value. Letting $N_{r}$ denote the number of arms in $A^{(r)}_{c}$ whose empirical average is larger than that of the optimal arm, we have:
$\mathbb{E}[N_{r}]$ $=\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\Pr\!\big{[}\hat{Q}^{\alpha,(r)}_{c,a_{1}}<\hat{Q}^{\alpha,(r)}_{c,a_{i}}\big{]}$
$\leq\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\exp\!\Big{(}-\frac{1}{2}harmonic(t_{r_{1}},t_{r_{i}})\Delta_{c,a_{i}}^{2}\Big{)}$
$\leq\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\exp\!\Big{(}-\Delta_{c,a_{i}}^{2}\cdot\tfrac{T}{2|\mathcal{P}|\bar{N}_{i}\log_{2}|\mathcal{A}|}\Big{)}$
$\leq\;|\mathcal{A}^{(r)}_{c}|\max_{i\in\mathcal{A}^{(r)}_{c}}\exp\!\Big{(}-\Delta_{c,a_{i}}^{2}\cdot\tfrac{T}{2|\mathcal{P}|\bar{N}_{max}\log_{2}|\mathcal{A}|}\Big{)}$
$\leq\;|\mathcal{A}^{(r)}_{c}|\exp\!\Big{(}-\tfrac{T}{2|\mathcal{P}|H_{1}^{c}R}\Big{)}$
[…]
$\Pr\!\Big{[}N_{r}>\tfrac{1}{2}|\mathcal{A}^{(r)}_{c}|\Big{]}\;\leq\;2\,\mathbb{E}[N_{r}]/|\mathcal{A}^{(r)}_{c}|\;\leq\;2\exp\!\Big{(}-\tfrac{T}{2|\mathcal{P}|H_{1}^{c}R}\Big{)}$
and the lemma follows. ∎
Lemma 2.
The probability that the best arm under context $c$ is eliminated from context $c$ on round $r$ is at most
$3\,\exp\!\Big{(}-\tfrac{T}{8|\mathcal{C}|H_{2}^{c}R}\Big{)}$
Proof.
The proof follows directly (Karnin, Koren, and Somekh 2013) Lemma 4.3. The only thing to recognize is that:
$\mathbb{E}[N_{r}]$ $=\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\Pr\!\big{[}\hat{Q}^{\alpha,(r)}_{c,a_{1}}<\hat{Q}^{\alpha,(r)}_{c,a_{i}}\big{]}$
$\leq\sum_{a_{i}\in\mathcal{A}^{(r)}_{c}}\exp\!\Big{(}-\Delta_{c,a_{i}}^{2}\cdot\tfrac{2^{r}T}{8|\mathcal{C}||\mathcal{A}|\bar{N}_{i}\log_{2}|\mathcal{A}|}\Big{)}$
Proof of Theorem 1.
The best arm needs to survive for all $R$ rounds and under all contexts $\mathcal{C}$. Therefore, from the Lemma 1:
$\sum_{r=1}^{R}\sum_{c}2\exp\!\Big{(}-\tfrac{T}{2|\mathcal{P}|H_{1}^{c}R}\Big{)}\;\leq\;3|\mathcal{C}|R\exp\!\Big{(}-\tfrac{T}{2|\mathcal{P}|H_{1}R}\Big{)}$
From the Lemma 2:
$\sum_{r=1}^{R}\sum_{c}3\exp\!\Big{(}-\tfrac{T}{8|\mathcal{C}|H_{2}^{c}R}\Big{)}\;\leq\;3|\mathcal{C}|R\exp\!\Big{(}-\tfrac{T}{8|\mathcal{C}|H_{2}R}\Big{)}$
Combining both:
$3|\mathcal{C}|R\exp\!\Big{(}-\tfrac{T}{min(2|\mathcal{P}|H_{1},8|\mathcal{C}|H_{2})R}\Big{)}$
which gives the theorem. ∎
Proposition 2 (Inference-Agnostic Optimality).
The Inference-Agnostic prompt-optimization policy remains optimal under linear transformation of $R_{x}^{\textsc{IA}}(c,a)$, that is, $kR_{x}^{\textsc{IA}}(c,a),k>0$ and an optimal policy can be recovered trivially from $Q$-function under affine transformation:
$Q^{AF}(c,a):=\mathbb{E}_{x\sim\mathcal{X}}\bigl{[}aR_{x}^{IA}(c,a)+b\bigr{]}=kQ^{IA}(c,a)+b.\quad$
Proof.
Follows directly from Jensen’s inequality. ∎
Appendix B
Synthetic-Bernoulli Environment.
[…]
A pull of $N\leq N_{\max}$ for prompt $p$ on example $x$ generates i.i.d. Bernoulli outcomes $\{c_{i}\}_{i=1}^{N}$ where $\Pr(c_{i}=1)=q_{p}(x)$, and each completion incurs a per-sample cost $k_{p}$. The result is an array $\bigl{[}\;c_{i},\;k_{p}\;\bigr{]}_{i=1}^{N}$.
Majority vote (MV) sets $M=1$ if $\sum_{i}c_{i}>N/2$, $M=0$ if $\sum_{i}c_{i}<N/2$, and assigns $M=0.5$ (by fair coin) in the case of a tie ($N$ even, $\sum_{i}c_{i}=N/2$).
The utility for cost for context $c\in\{\text{low},\text{mid},\text{high}\}$ is computed as
$u_{c}=w_{1}M+w_{2}(c)\sum_{i=1}^{N}k_{p},$
where $w_{1}=1$ and $w_{2}(c)\in\{0,\ -0.2,\ -1.0\}$ depending on the cost tier.
To instantiate the environment, we generate two prompt archetypes: deceiving prompts, which achieve high average accuracy but exhibit low $q_{p}(x)$ on hard queries, and all-rounders, which maintain moderate accuracy more uniformly across tiers. Per-prompt costs $k_{p}$ are sampled from a normal distribution with mean $0.02$ and variance $0.005$.
Synthetic-Categorical Environment.
We model $P=32$ prompts, each paired with $|\mathcal{X}|=512$ queries and $K=2$ positive objectives. For every $(p,x)$, there are $M$ categorical outcomes, each represented by a vector $o_{j}\in\mathbb{R}^{K}$. A pull of $N\leq N_{\max}(=32)$ for prompt $p$ on query $x$ generates $N$ i.i.d. outcome vectors, resulting in rows $o_{i,1},\,o_{i,2},\,k_{p}\;]$, where $k_{p}$ denotes the per-completion cost for prompt $p$.
Given a context $c$ with weights $w=(w_{1},w_{2},w_{\text{cost}})$, where $w_{1}+w_{2}=1$ and $w_{\text{cost}}\leq 0$, the Best-of-$N$ utility is defined as
$u_{c}=\max_{1\leq i\leq N}\left(w_{1}o_{i,1}+w_{2}o_{i,2}\right)+w_{\text{cost}}\,N\,k_{p}.$
To construct the environment, outcome vectors are sampled from $\{-4,\ldots,4\}^{2}$. We instantiate two prompt archetypes: HMLV (high mean, low variance; excels at N=1N{=}1) and LMHV (lower mean, high variance; benefits from larger $N$), each specializing in one objective. For each $(p,x)$, we add small per-query noise to the categorical outcome probabilities, introduce a mild train-to-test shift by perturbing these probabilities, sample per-prompt costs $k_{p}\in[0.02,\,0.1]$, and draw context weights from a grid satisfying $w_{1}+w_{2}=1$ with $w_{\text{cost}}\in\{-0.1,\ -0.5,\ -1.0\}$.
MATH Environment.
We select $316$ integer-answer problems from the MATH dataset [^2] https://huggingface.co/datasets/HuggingFaceH4/MATH-500. A set of $25$ prompt templates is authored using ChatGPT-o3. For each $(\text{prompt},\,\text{problem})$ pair, we sample $128$ responses from Llama-3.3-70B-Instruct at temperature $T=0.7$, parsing each completion to its final integer answer.
The dataset is then processed as follows:
  1. 1.
    For each problem, retain the global top-4 answers and group all other answers into a single other bucket ($C=5$ categories in total).
  2. 2.
    Compute per-prompt costs as the normalized average token length of its responses.
This yields a categorical environment (analogous to the Synthetic-Categorical setting) with $P=25$, $N_{\max}=32$, a uniform context prior $c\in\{\text{low},\,\text{mid},\,\text{high}\}$, and cost coefficients {0, $\{0,\,0.2,\,1$.0\}. Utility is evaluated via majority vote.
CommonsenseQA Environment.
We randomly sample 1,5001{,}500 multiple-choice questions from the CommonsenseQA corpus [^3] https://huggingface.co/datasets/tau/commonsense˙qa, and author $48$ prompt templates using ChatGPT-o3. For each $(\text{prompt},\,\text{question})$ pair, we query Llama-3.3-70B-Instruct at temperature $T=1.1$, collecting $128$ JSON-constrained answers (one of “Option A”–“Option E”). Each prompt is assigned a constant cost $k_{p}=0.01$.
The resulting data is used to construct a categorical environment (in analogy to the Synthetic-Categorical setting) with $P=48$, $N_{\max}=32$, a uniform context prior, and cost coefficients {0, $\{0,\,0.2,\,1$.0\}.
Helpful–Harmless Environment.
We filter the HH-RLHF conversations [^4] https://huggingface.co/datasets/Anthropic/hh-rlhf to the 1,3551{,}355 examples containing a single user query and a single assistant response. Using ChatGPT-o3, we craft $20$ prompt templates. For each $(\text{prompt},\,\text{query})$ pair, we sample $128$ continuations from Llama-3.3-70B-Instruct at temperature $T=0.7$. Each continuation is scored by separate public reward models (Yang et al. 2024) for helpfulness [^5] Ray2333/gpt2-large-helpful-reward˙model and harmlessness [^6] Ray2333/gpt2-large-harmless-reward˙model, with scores normalized to [1,1][-1,1].
The two reward scores are then binned on a $0.5$-spaced grid, producing a categorical distribution per $(\text{prompt},\,\text{query})$; per-prompt costs are computed as the average token length. This data defines a categorical environment with $P=20$, $N_{\max}=32$, a uniform context prior over weight triples $(w_{h},\,w_{s},\,w_{\text{cost}})$ with $w_{h}+w_{s}=1$ and $w_{\text{cost}}\in\{-0.1,\,-0.5,\,-1.0\}$.
Summarization Environment.
We randomly sample 1,2011{,}201 Reddit posts from the Summarize-from-Feedback corpus [^7] https://huggingface.co/datasets/openai/summarize˙from˙feedback and design $20$ summarization prompt templates using ChatGPT-o3. For each $(\text{prompt},\,\text{post})$ pair, we query Llama-3.3-70B-Instruct at temperature $T=0.7$ and collect $128$ candidate summaries.
Each summary is scored by two publicly available reward models: Preference [^8] OpenAssistant/reward-model-deberta-v3-large-v2 and Faithful [^9] CogComp/bart-faithful-summary-detector, with raw scores normalized to [1,1][-1,1]. We then bin each dimension in steps of $0.5$, producing a categorical distribution over the two reward dimensions, and compute per-prompt costs from average token length.
This data defines a categorical environment with $P=20$, $N_{\max}=32$, and a uniform context prior over weight triples $(w_{h},\,w_{s},\,w_{\text{cost}})$ where $w_{h}+w_{s}=1$ and $w_{\text{cost}}\in\{-0.1,\,-0.5,\,-1.0\}$.
Note: All prompts are available under the prompts folder of the code base.
Appendix C
Top-$K$ screening.
For the screening variant, we fixed $K=4$ candidates after screening and swept the burn-in fraction $\rho\in\{0.05,0.10,0.20,0.30,0.40\}$, which allocates a $\rho$-portion of the budget to obtain initial estimates before trimming. Parameter sweep protocol matched the baselines. We selected $\rho=0.20$ for reporting, as it achieved the best overall performance while remaining robust across datasets and inference regimes 2.
UCB.
We tuned the exploration constant over $c\in\{0.1,0.5,1.0,2.0,4.0,8.0\}$ under the same budgets and using 20% of the data per environment; identical seeds across settings; 10,00010{,}000 test contexts). The agent ranks arms by the standard UCB index
$\mathrm{UCB}_{i}(t)\;=\;\hat{\mu}_{i}(t)\;+\;c\,\sqrt{\frac{\ln t}{n_{i}(t)}},$
where $\hat{\mu}_{i}(t)$ is the empirical mean utility of arm $i$, $n_{i}(t)$ its pull count, and $t$ the total pulls. We selected $c=0.1$ for reporting, as it achieved the best overall performance while remaining robust across datasets and inference regimes 3.
$\epsilon$-greedy.
We swept $\varepsilon\in\{0.50,0.75,0.80,0.85,0.90,0.95\}$ separately for each dataset and inference regime (MV, BoN). For every $\varepsilon$, agents were trained under budgets $T\in\{3\text{K},5\text{K},10\text{K},20\text{K},30\text{K},40\text{K}\}$, using 20% of the data per environment with deterministic reseeding; evaluation used 10,00010{,}000 test contexts per environment. We selected $\epsilon=0.15$ for reporting, as it achieved the best overall performance while remaining robust across datasets and inference regimes 4.
Param $\times$ T $H$ Summarization SC SB MATH CQA
$\rho$=0.05, $T=3000$ 0.40 $\pm$ 0.00 0.20 $\pm$ 0.00 2.77 $\pm$ 0.02 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
$\rho$=0.05, $T=5000$ 0.40 $\pm$ 0.00 0.22 $\pm$ 0.00 2.83 $\pm$ 0.02 0.83 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.05, $T=10000$ 0.42 $\pm$ 0.00 0.21 $\pm$ 0.00 2.83 $\pm$ 0.02 0.85 $\pm$ 0.01 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.05, $T=20000$ 0.43 $\pm$ 0.00 0.22 $\pm$ 0.00 2.87 $\pm$ 0.02 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.05, $T=30000$ 0.44 $\pm$ 0.00 0.23 $\pm$ 0.00 2.88 $\pm$ 0.01 0.84 $\pm$ 0.00 0.82 $\pm$ 0.01 0.77 $\pm$ 0.00
$\rho$=0.05, $T=40000$ 0.43 $\pm$ 0.00 0.23 $\pm$ 0.00 2.87 $\pm$ 0.02 0.84 $\pm$ 0.00 0.81 $\pm$ 0.00 0.77 $\pm$ 0.00
$\rho$=0.10, $T=3000$ 0.41 $\pm$ 0.00 0.21 $\pm$ 0.00 2.79 $\pm$ 0.02 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.10, $T=5000$ 0.41 $\pm$ 0.00 0.22 $\pm$ 0.00 2.84 $\pm$ 0.02 0.84 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.10, $T=10000$ 0.42 $\pm$ 0.00 0.21 $\pm$ 0.00 2.86 $\pm$ 0.02 0.84 $\pm$ 0.00 0.81 $\pm$ 0.01 0.77 $\pm$ 0.00
$\rho$=0.10, $T=20000$ 0.43 $\pm$ 0.00 0.23 $\pm$ 0.00 2.88 $\pm$ 0.02 0.84 $\pm$ 0.00 0.81 $\pm$ 0.01 0.77 $\pm$ 0.00
$\rho$=0.10, $T=30000$ 0.44 $\pm$ 0.00 0.23 $\pm$ 0.00 2.89 $\pm$ 0.02 0.84 $\pm$ 0.00 0.81 $\pm$ 0.01 0.77 $\pm$ 0.01
$\rho$=0.10, $T=40000$ 0.44 $\pm$ 0.00 0.23 $\pm$ 0.00 2.88 $\pm$ 0.02 0.84 $\pm$ 0.00 0.82 $\pm$ 0.00 0.77 $\pm$ 0.00
$\rho$=0.20, $T=3000$ 0.41 $\pm$ 0.00 0.20 $\pm$ 0.00 2.77 $\pm$ 0.02 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.20, $T=5000$ 0.41 $\pm$ 0.00 0.22 $\pm$ 0.00 2.84 $\pm$ 0.02 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.20, $T=10000$ 0.43 $\pm$ 0.00 0.22 $\pm$ 0.00 2.85 $\pm$ 0.02 0.83 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.20, $T=20000$ 0.43 $\pm$ 0.00 0.23 $\pm$ 0.00 2.87 $\pm$ 0.01 0.84 $\pm$ 0.00 0.81 $\pm$ 0.01 0.77 $\pm$ 0.00
$\rho$=0.20, $T=30000$ 0.44 $\pm$ 0.00 0.23 $\pm$ 0.00 2.89 $\pm$ 0.02 0.84 $\pm$ 0.00 0.82 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.20, $T=40000$ 0.44 $\pm$ 0.00 0.22 $\pm$ 0.00 2.88 $\pm$ 0.02 0.84 $\pm$ 0.00 0.82 $\pm$ 0.01 0.77 $\pm$ 0.00
$\rho$=0.30, $T=3000$ 0.41 $\pm$ 0.00 0.21 $\pm$ 0.00 2.81 $\pm$ 0.02 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.30, $T=5000$ 0.41 $\pm$ 0.00 0.22 $\pm$ 0.00 2.85 $\pm$ 0.01 0.83 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.30, $T=10000$ 0.42 $\pm$ 0.00 0.22 $\pm$ 0.00 2.85 $\pm$ 0.02 0.84 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.30, $T=20000$ 0.43 $\pm$ 0.00 0.22 $\pm$ 0.00 2.88 $\pm$ 0.01 0.83 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.30, $T=30000$ 0.44 $\pm$ 0.00 0.22 $\pm$ 0.00 2.88 $\pm$ 0.01 0.84 $\pm$ 0.00 0.82 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.30, $T=40000$ 0.44 $\pm$ 0.00 0.23 $\pm$ 0.00 2.89 $\pm$ 0.01 0.84 $\pm$ 0.00 0.82 $\pm$ 0.01 0.77 $\pm$ 0.00
$\rho$=0.40, $T=3000$ 0.41 $\pm$ 0.00 0.20 $\pm$ 0.00 2.80 $\pm$ 0.01 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.40, $T=5000$ 0.42 $\pm$ 0.00 0.20 $\pm$ 0.00 2.80 $\pm$ 0.02 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
$\rho$=0.40, $T=10000$ 0.43 $\pm$ 0.00 0.22 $\pm$ 0.00 2.85 $\pm$ 0.01 0.83 $\pm$ 0.00 0.81 $\pm$ 0.01 0.77 $\pm$ 0.00
$\rho$=0.40, $T=20000$ 0.43 $\pm$ 0.00 0.22 $\pm$ 0.00 2.87 $\pm$ 0.02 0.83 $\pm$ 0.00 0.81 $\pm$ 0.01 0.77 $\pm$ 0.00
$\rho$=0.40, $T=30000$ 0.44 $\pm$ 0.00 0.22 $\pm$ 0.00 2.88 $\pm$ 0.01 0.83 $\pm$ 0.00 0.81 $\pm$ 0.01 0.77 $\pm$ 0.00
$\rho$=0.40, $T=40000$ 0.44 $\pm$ 0.00 0.23 $\pm$ 0.00 2.89 $\pm$ 0.01 0.84 $\pm$ 0.00 0.82 $\pm$ 0.01 0.77 $\pm$ 0.00
Table 2: PSST+K4: mean $\pm$ SEM across datasets (rows are param, $\rho$ and $T$).
Param $\times$ T $H$ Summarization SC SB MATH CQA
c=0.1, $T=3000$ 0.38 $\pm$ 0.00 0.19 $\pm$ 0.01 2.83 $\pm$ 0.02 0.82 $\pm$ 0.00 0.78 $\pm$ 0.01 0.75 $\pm$ 0.00
c=0.1, $T=5000$ 0.39 $\pm$ 0.00 0.21 $\pm$ 0.00 2.88 $\pm$ 0.01 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
c=0.1, $T=10000$ 0.41 $\pm$ 0.00 0.23 $\pm$ 0.00 2.94 $\pm$ 0.01 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
c=0.1, $T=20000$ 0.43 $\pm$ 0.00 0.25 $\pm$ 0.00 2.98 $\pm$ 0.01 0.86 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
c=0.1, $T=30000$ 0.43 $\pm$ 0.00 0.25 $\pm$ 0.00 2.99 $\pm$ 0.01 0.88 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.01
c=0.1, $T=40000$ 0.44 $\pm$ 0.00 0.25 $\pm$ 0.00 3.00 $\pm$ 0.01 0.89 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
c=0.5, $T=3000$ 0.37 $\pm$ 0.00 0.19 $\pm$ 0.01 2.85 $\pm$ 0.02 0.82 $\pm$ 0.00 0.78 $\pm$ 0.01 0.75 $\pm$ 0.00
c=0.5, $T=5000$ 0.38 $\pm$ 0.00 0.20 $\pm$ 0.00 2.90 $\pm$ 0.01 0.82 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
c=0.5, $T=10000$ 0.41 $\pm$ 0.00 0.23 $\pm$ 0.00 2.94 $\pm$ 0.01 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
c=0.5, $T=20000$ 0.43 $\pm$ 0.00 0.24 $\pm$ 0.00 2.98 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
c=0.5, $T=30000$ 0.43 $\pm$ 0.00 0.24 $\pm$ 0.00 3.00 $\pm$ 0.01 0.86 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
c=0.5, $T=40000$ 0.44 $\pm$ 0.00 0.25 $\pm$ 0.00 3.00 $\pm$ 0.01 0.88 $\pm$ 0.00 0.81 $\pm$ 0.01 0.75 $\pm$ 0.00
c=1.0, $T=3000$ 0.37 $\pm$ 0.00 0.19 $\pm$ 0.01 2.88 $\pm$ 0.02 0.82 $\pm$ 0.00 0.78 $\pm$ 0.01 0.75 $\pm$ 0.00
c=1.0, $T=5000$ 0.37 $\pm$ 0.00 0.19 $\pm$ 0.01 2.91 $\pm$ 0.02 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
c=1.0, $T=10000$ 0.41 $\pm$ 0.00 0.22 $\pm$ 0.00 2.94 $\pm$ 0.01 0.83 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
c=1.0, $T=20000$ 0.42 $\pm$ 0.00 0.24 $\pm$ 0.00 2.98 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
c=1.0, $T=30000$ 0.43 $\pm$ 0.00 0.24 $\pm$ 0.00 3.00 $\pm$ 0.01 0.86 $\pm$ 0.01 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
c=1.0, $T=40000$ 0.43 $\pm$ 0.00 0.25 $\pm$ 0.00 3.00 $\pm$ 0.01 0.87 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.01
c=2.0, $T=3000$ 0.37 $\pm$ 0.00 0.18 $\pm$ 0.01 2.86 $\pm$ 0.02 0.82 $\pm$ 0.00 0.78 $\pm$ 0.01 0.75 $\pm$ 0.00
c=2.0, $T=5000$ 0.38 $\pm$ 0.00 0.19 $\pm$ 0.01 2.93 $\pm$ 0.01 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.76 $\pm$ 0.00
c=2.0, $T=10000$ 0.40 $\pm$ 0.00 0.23 $\pm$ 0.00 2.94 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
c=2.0, $T=20000$ 0.42 $\pm$ 0.00 0.24 $\pm$ 0.00 2.98 $\pm$ 0.01 0.85 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
c=2.0, $T=30000$ 0.42 $\pm$ 0.00 0.24 $\pm$ 0.00 2.99 $\pm$ 0.01 0.86 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
c=2.0, $T=40000$ 0.43 $\pm$ 0.00 0.25 $\pm$ 0.00 2.99 $\pm$ 0.01 0.88 $\pm$ 0.00 0.81 $\pm$ 0.01 0.75 $\pm$ 0.00
c=4.0, $T=3000$ 0.37 $\pm$ 0.00 0.18 $\pm$ 0.01 2.85 $\pm$ 0.02 0.82 $\pm$ 0.00 0.78 $\pm$ 0.01 0.75 $\pm$ 0.00
c=4.0, $T=5000$ 0.37 $\pm$ 0.00 0.18 $\pm$ 0.00 2.91 $\pm$ 0.01 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.76 $\pm$ 0.00
c=4.0, $T=10000$ 0.41 $\pm$ 0.00 0.22 $\pm$ 0.00 2.94 $\pm$ 0.01 0.84 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
c=4.0, $T=20000$ 0.42 $\pm$ 0.00 0.24 $\pm$ 0.00 2.97 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
c=4.0, $T=30000$ 0.42 $\pm$ 0.00 0.24 $\pm$ 0.00 2.99 $\pm$ 0.01 0.86 $\pm$ 0.00 0.81 $\pm$ 0.01 0.75 $\pm$ 0.00
c=4.0, $T=40000$ 0.43 $\pm$ 0.00 0.25 $\pm$ 0.00 3.00 $\pm$ 0.01 0.87 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.00
c=8.0, $T=3000$ 0.37 $\pm$ 0.00 0.18 $\pm$ 0.01 2.86 $\pm$ 0.02 0.82 $\pm$ 0.00 0.78 $\pm$ 0.01 0.75 $\pm$ 0.00
c=8.0, $T=5000$ 0.38 $\pm$ 0.00 0.19 $\pm$ 0.01 2.90 $\pm$ 0.02 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.76 $\pm$ 0.00
c=8.0, $T=10000$ 0.40 $\pm$ 0.00 0.22 $\pm$ 0.00 2.92 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
c=8.0, $T=20000$ 0.42 $\pm$ 0.00 0.24 $\pm$ 0.00 2.97 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
c=8.0, $T=30000$ 0.43 $\pm$ 0.00 0.24 $\pm$ 0.00 2.98 $\pm$ 0.01 0.86 $\pm$ 0.00 0.81 $\pm$ 0.01 0.75 $\pm$ 0.00
c=8.0, $T=40000$ 0.43 $\pm$ 0.00 0.25 $\pm$ 0.00 2.99 $\pm$ 0.01 0.87 $\pm$ 0.00 0.81 $\pm$ 0.01 0.76 $\pm$ 0.01
Table 3: UCB: mean $\pm$ SEM across datasets (rows are param, $T$).
Param $\times$ T $H$ Summarization SC SB MATH CQA
e=0.50, $T=3000$ 0.37 $\pm$ 0.00 0.17 $\pm$ 0.01 2.78 $\pm$ 0.02 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.50, $T=5000$ 0.39 $\pm$ 0.00 0.20 $\pm$ 0.01 2.82 $\pm$ 0.01 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.50, $T=10000$ 0.41 $\pm$ 0.00 0.21 $\pm$ 0.00 2.90 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.50, $T=20000$ 0.42 $\pm$ 0.00 0.23 $\pm$ 0.00 2.94 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.01 0.74 $\pm$ 0.00
e=0.50, $T=30000$ 0.43 $\pm$ 0.00 0.24 $\pm$ 0.00 2.97 $\pm$ 0.01 0.85 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.50, $T=40000$ 0.43 $\pm$ 0.00 0.25 $\pm$ 0.00 2.98 $\pm$ 0.01 0.85 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.01
e=0.55, $T=3000$ 0.38 $\pm$ 0.00 0.16 $\pm$ 0.01 2.75 $\pm$ 0.03 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.75, $T=5000$ 0.39 $\pm$ 0.00 0.17 $\pm$ 0.01 2.86 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.75, $T=10000$ 0.40 $\pm$ 0.00 0.20 $\pm$ 0.01 2.91 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.75, $T=20000$ 0.42 $\pm$ 0.00 0.23 $\pm$ 0.00 2.95 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.00 0.74 $\pm$ 0.00
e=0.75, $T=30000$ 0.43 $\pm$ 0.00 0.23 $\pm$ 0.00 2.97 $\pm$ 0.01 0.85 $\pm$ 0.00 0.81 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.75, $T=40000$ 0.43 $\pm$ 0.00 0.24 $\pm$ 0.00 2.96 $\pm$ 0.01 0.85 $\pm$ 0.00 0.80 $\pm$ 0.01 0.74 $\pm$ 0.00
e=0.80, $T=3000$ 0.38 $\pm$ 0.00 0.18 $\pm$ 0.01 2.83 $\pm$ 0.02 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.80, $T=5000$ 0.39 $\pm$ 0.00 0.19 $\pm$ 0.00 2.86 $\pm$ 0.02 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.80, $T=10000$ 0.40 $\pm$ 0.00 0.19 $\pm$ 0.01 2.91 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.80, $T=20000$ 0.42 $\pm$ 0.00 0.23 $\pm$ 0.00 2.94 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.80, $T=30000$ 0.41 $\pm$ 0.00 0.23 $\pm$ 0.00 2.96 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.80, $T=40000$ 0.43 $\pm$ 0.00 0.24 $\pm$ 0.00 2.98 $\pm$ 0.01 0.85 $\pm$ 0.00 0.80 $\pm$ 0.01 0.74 $\pm$ 0.00
e=0.85, $T=3000$ 0.38 $\pm$ 0.00 0.16 $\pm$ 0.01 2.72 $\pm$ 0.04 0.83 $\pm$ 0.00 0.78 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.85, $T=5000$ 0.38 $\pm$ 0.00 0.17 $\pm$ 0.01 2.87 $\pm$ 0.01 0.83 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.85, $T=10000$ 0.40 $\pm$ 0.00 0.20 $\pm$ 0.00 2.90 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.85, $T=20000$ 0.41 $\pm$ 0.00 0.22 $\pm$ 0.00 2.95 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.85, $T=30000$ 0.42 $\pm$ 0.00 0.23 $\pm$ 0.01 2.95 $\pm$ 0.01 0.85 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.85, $T=40000$ 0.42 $\pm$ 0.00 0.24 $\pm$ 0.00 2.97 $\pm$ 0.01 0.85 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.01
e=0.90, $T=3000$ 0.37 $\pm$ 0.00 0.17 $\pm$ 0.01 2.81 $\pm$ 0.03 0.82 $\pm$ 0.00 0.78 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.90, $T=5000$ 0.38 $\pm$ 0.00 0.17 $\pm$ 0.01 2.87 $\pm$ 0.02 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.90, $T=10000$ 0.40 $\pm$ 0.00 0.19 $\pm$ 0.01 2.90 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.90, $T=20000$ 0.41 $\pm$ 0.00 0.22 $\pm$ 0.00 2.94 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.90, $T=30000$ 0.42 $\pm$ 0.00 0.23 $\pm$ 0.00 2.95 $\pm$ 0.01 0.85 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.90, $T=40000$ 0.42 $\pm$ 0.00 0.24 $\pm$ 0.00 2.97 $\pm$ 0.01 0.85 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.95, $T=3000$ 0.37 $\pm$ 0.00 0.17 $\pm$ 0.01 2.76 $\pm$ 0.03 0.82 $\pm$ 0.00 0.78 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.95, $T=5000$ 0.38 $\pm$ 0.00 0.17 $\pm$ 0.01 2.86 $\pm$ 0.01 0.83 $\pm$ 0.00 0.79 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.95, $T=10000$ 0.39 $\pm$ 0.00 0.19 $\pm$ 0.01 2.92 $\pm$ 0.01 0.84 $\pm$ 0.00 0.80 $\pm$ 0.01 0.76 $\pm$ 0.00
e=0.95, $T=20000$ 0.41 $\pm$ 0.00 0.21 $\pm$ 0.00 2.95 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.95, $T=30000$ 0.42 $\pm$ 0.00 0.22 $\pm$ 0.00 2.95 $\pm$ 0.01 0.85 $\pm$ 0.00 0.80 $\pm$ 0.01 0.75 $\pm$ 0.00
e=0.95, $T=40000$ 0.41 $\pm$ 0.00 0.23 $\pm$ 0.00 2.95 $\pm$ 0.01 0.84 $\pm$ 0.00 0.79 $\pm$ 0.00 0.75 $\pm$ 0.00
Table 4: $\epsilon$-greedy: mean $\pm$ SEM across datasets (rows are param, $T$).
Appendix D
Statistical testing.
For each dataset and budget $T$, we perform all pairwise algorithm comparisons using per-environment utilities as paired samples (identical train/test splits via deterministic reseeding). Our default test is the two-sided Wilcoxon signed-rank test, which we apply to the aligned vectors after removing non-finite values and dropping exact ties (zero_method=wilcox, mode=auto); pairs with fewer than two effective samples are skipped. When requested, we also report the paired sign test (binomial test on the sign of differences) after removing ties. To control multiplicity within each $(\text{dataset},T)$ grid, we use Holm–Bonferroni adjustment by default (with options for Benjamini–Hochberg FDR or no correction). We declare a winner if the adjusted $p<\alpha=0.05$; the direction is determined by the sign of the median difference $\operatorname{median}(x-y)$. In case of unequal environment counts across algorithms, samples are truncated to the minimum length to preserve pairing. Figures visualize the outcome matrix with entries in $\{-1,0,+1\}$ indicating row-algorithm loss, non-significance, or win against the column algorithm, respectively.
All the results are shown in Figs 56789,  10. Across all six datasets, we observe that PSST and the Top-K screening heuristic consistently outperform competing methods across most budget settings, with statistical significance.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Pairwise wins for Commonsense QA (MV) across six budgets ($T$ in order: 3000, 5000, 10000, 20000, 30000, 40000).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Pairwise wins for Helpful-Harmless (BoN) across six budgets ($T$ in order: 3000, 5000, 10000, 20000, 30000, 40000).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Pairwise wins for MATH (MV) across six budgets ($T$ in order: 3000, 5000, 10000, 20000, 30000, 40000).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 8: Pairwise wins for Synthetic Bernoulli (MV) across six budgets ($T$ in order: 3000, 5000, 10000, 20000, 30000, 40000).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 9: Pairwise wins for Synthetic Categorical (BoN) across six budgets ($T$ in order: 3000, 5000, 10000, 20000, 30000, 40000).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: Pairwise wins for Summarization (BoN) across six budgets ($T$ in order: 3000, 5000, 10000, 20000, 30000, 40000).